# Predicting the Geolocation of Tweets Using transformer models on Customized Data

### v5

@@ 2 Related works
The most straightforward approach is to analyze the natural languages present in social media posts. There are multiple projects processing documents or the summary of user profile tweets to predict the author geolocation [28] [37] [11]. These works apply geographical grids to divide the Earth’s surface into different-sized regions such that highly populated areas are split into smaller pieces while more sparse population areas end up as large grid cells. Early efforts were mainly focused on mining indicative information from users’ posting content relying on location indicative words (LIWs) that can link users to their home locations, based on various NLP techniques (e.g., topic models and statistic models) [46]. For example, Rahimi extracts bag-of-words features from user posts [26]; Wing and Baldridge estimate the word distributions for different regions [36]. Other word-centric works [32] [23] [42] also focus on filtering LIWs from the text and using the gazetteers including not only city/country names but also dialect terms to resolve the geographical indexes. While such methodologies suppose the predefined set of LIWs and their geographical coordinates, Rahimi [25] proposes a neural network encoding of such words and phrases to the continuous two-dimensional space. In this light, we aimed to develop another semi-supervised method based on the state-of-the-art models for NLP with no need for LIWs datasets or geographic gazetteers.
@@
Despite short in-length textual content, a tweet is accompanied by rich context like timestamps and geo-tags associated with it, as well as various attributes of its author profile. For instance, when sending out a tweet, it is automatically accompanied by a posting timestamp. Moreover, in the era of widespread GPS-enabled devices such as smartphones and tablets, users have the option to append geo-tags to their tweets, indicating their current locations [44]. In the case of Twitter, the spatial context is represented as machine-generated geographic object descriptions with countries, cities, place types, and full names attributed by Twitter to the user-defined geolocation coordinates.
@@
In terms of the tweet’s author profile metadata, Schulz in the work [32], incorporates users’ self-declared home locations, personal websites, and currently outdated timezone attributes, as potential tweet location indicators. For instance, the personal websites were processed to obtain the country code and the website’s geocoded IP address as spatial indicators. While the extracted features offer valuable insights into a user’s home country, they also lead to a noticeable increase in tweet processing time. Given the imperative of maintaining a swift processing pace, especially when dealing with an overall Twitter feed generation rate of 6,000 tweets per second, the research focus shifted to utilizing readily available fields of the tweet’s author profile. In particular, textual fields such as users’ home locations, profile bios, names, and screen names were used as model input features for the tweet location prediction.
@@
Importantly, Priedhorsky’s methodology was aligned with the present work as it utilized GMMs for geolocation prediction instead of simple coordinates prediction [24]. Another subsequent study, which employed a hybrid approach, also estimated geolocation through GMMs rather than coordinates [2]. This study used the text and network features jointly as predictors for the final estimation. The performance metrics used in both works were adopted to evaluate the proposed probabilistic models as described in Section B.2.
@@
Most of the previously mentioned works use various approaches based on neural networks for NLP, but none of them has applied transformer models for the estimation of geolocation. The relatively new BERT model (released in 2018 [7]) is aimed at classification tasks and some works have already used it for country/city names predictions.
The base BERT model is intended to be finetuned on a downstream task which, in this case, was a regression to geographical coordinates (and GMM parameters such as weights and covariance) output as a type of sentence classification task. In general, BERT models have been pretrained using Masked Language Modeling (MLM), which enables the model to learn bidirectional representations of sentences, and Next Sentence Prediction (NSP) to capture the relationships between sentences. Since BERT’s layers are hierarchical, early BERT layers learn more generic linguistic patterns, such as differences between multiple languages and their general sentence structures (in the case of the BERT multilingual model). While the later layers learn more task-specific patterns, in this case—the associations between points on the world map and specific terms referred to as LIWs, as well as text constructions and linguistic patterns commonly used in certain geographical areas.
@@
Nevertheless, the global dataset necessitated the use of the multilingual BERT base model (’bert-base-multilingual-cased’), which resulted in higher error distance metrics for the per-tweet and per-user geolocation prediction, as expected. A critical challenge addressed is the representation of multiple languages, since the ability to process multiple languages is a key factor in dividing text inputs into geospatial regions. To tackle this issue, a multilingual BERT model, pretrained on the 104 largest Wikipedia languages, was utilized as a starting point for finetuning on a global dataset with geolocation labels.
@@ 3 Materials and Methods
Step 2: importing default tokenizer and base BERT [7] model;
@@
The hyper-parameters tuning stage has covered such parameters as the type of the scheduler, minimal and maximal learning rates for the chosen scheduler, and the number of epochs. For the current task, the cosine type was chosen from the set of linear, cyclic, step, and plateau schedulers. Experiments have shown that the optimal learning rate reduction range starts at 1e-5 and ends at 1e-6 at the end of the last training epoch. The entering number of epochs was reduced from 5 to 3 for the dataset consisting of 2.7 million training samples, which are grouped into batches of 10 to 16 tweets in the model data loader. This decision was made as a result of the observed lack of significant error distance reduction after the third epoch for both training and test (development set) metrics. By default, the development set of 300,000 samples was evaluated without a gradient propagation at the end of each epoch. The described parameters were held constant across all proposed models, with variations only occurring in the loss function type, the number of outcomes (prediction points), covariance type, and the combination of text features.
@@ 3.1 Data preprocessing
Worth mentioning that the word-centric techniques, which use geocoders to determine the place that is being referred to, are based on the geographic gazetteers [32]. The gazetteer typically includes entries for cities, towns, villages, landmarks, natural features, and other geographic entities, along with information such as their location, coordinates, elevation, etc. The challenge in this work was to find a substitute for the accurate knowledge base of geographical references (e.g., gazetteer terms).
@@
Considering the challenges of computational complexity and data availability, we chose to concentrate solely on user profile textual information (name, screen name, description, and location of the Twitter user profile), as shown in Table 1. This allowed us to effectively achieve our research objectives of predicting geolocation for the arbitrary tweets within our available resources and datasets. Worth mentioning, that the user’s home location wasn’t retrieved from the profile but rather sampled from the user’s tweets subset during evaluation.
@@ 4 Discussion and Conclusion
The proposed model trained on Twitter data could be used to predict the location of any Twitter user by processing the publicly available information posted online. While profile information provided by Twitter users is generally considered public, the combination or analysis of such data can have privacy implications. Tweets, being public by default, can also be used for analysis and publication. However, it’s worth noting that even seemingly innocuous information, when combined with other data points or analyzed in aggregate, can potentially reveal more about a person’s identity or habits. Therefore, while individual profile details may not be inherently sensitive, their combination or analysis might raise privacy concerns.
@@
As for the real-time user location monitoring, Yuan in the work [40] focused on tracking the movement of individuals or groups over time. Monitoring user geolocation in time can provide valuable insights for a variety of applications such as tracking disease outbreaks, analyzing urban mobility patterns, and providing location-based services. Such an analytical framework could be built on top of the proposed models, yet, in this study, only the commonly researched task of the user’s home location prediction was properly explored so far.
@@ 5 Acknowledgments
The authors acknowledge the Institute of Science and Technology (ISTA) for their material support and for granting access to the Twitter database archive, which was essential for the research. We also extend our appreciation to ChatGPT, the language model developed by OpenAI, for its assistance in manuscript editing, coding documentation, and revising parts of the manuscript. The technical information and statements were formulated by the authors, and ChatGPT was used to transfer the text into more fluent English. The authors ensured that only the style and not the meaning were changed.

### v6

@@ 2 Related works
The most straightforward approach would be to process the natural languages present in social media posts. There are multiple projects processing documents or the summary of user profile tweets to predict the author geolocation [27] [36] [10]. These works apply geographical grids dividing the Earth’s globe surface into different-sized regions such that highly populated areas are split into smaller pieces while more sparse population areas end up as large grid cells. Early efforts were mainly focused on mining indicative information from users’ posting content relying on location indicative words (LIWs) that can link users to their home locations, based on various NLP techniques (e.g., topic models and statistic models) [45]. For example, Rahimi extracts bag-of-words features from user posts [25]; Wing and Baldridge estimate the word distributions for different regions [35]. Other word-centric works [31] [22] [41] also focus on filtering LIWs from the text and using the gazetteers including not only city/country names but also dialect terms to resolve the geographical indexes. While such methodologies suppose the predefined set of LIWs and their geographical coordinates, Rahimi [24] proposes a neural network encoding of such words and phrases to the continuous two-dimensional space. In this light, we aimed to develop another semi-supervised method based on the state-of-the-art models for NLP with no need for LIWs datasets or geographic gazetteers.
@@
Despite short in-length textual content, a tweet is accompanied by rich context like timestamps and geo-tags associated with it, and multiple attributes of its author profile. For instance, when sending out a tweet, it is automatically accompanied by a posting timestamp. Moreover, GPS-enabled devices such as smartphones and tablets have become widespread allowing users to easily append geo-tags to their tweets, indicating their current locations [43]. In the case of Twitter, the spatial context is represented as machine-generated geographic object descriptions with countries, cities, place types, and full names attributed by Twitter to the user-defined geolocation coordinates.
@@
In terms of the tweet’s author profile metadata, Schulz in the work [31], incorporates users’ self-declared home locations, personal websites, and currently outdated timezone attributes, as potential tweet location indicators. For instance, the personal websites were processed to obtain the spatial indicators such as country code and IP address. While the extracted features offer valuable insights into a user’s home country, they also lead to a noticeable increase in tweet processing time. Given the imperative of maintaining a swift processing pace, especially when dealing with an overall Twitter feed generation rate of 6,000 tweets per second, the research focus shifted to utilizing readily available fields of the tweet’s author profile. In particular, textual fields such as users’ home locations, profile bios, names, and screen names were used as model input features for the tweet location prediction.
@@
Importantly, Priedhorsky’s methodology was aligned with the present work as it utilized GMMs for geolocation prediction instead of simple coordinates prediction [23]. Another subsequent study, which employed a hybrid approach, also estimated geolocation through GMMs rather than coordinates [2]. This study’s final estimation was based on the text and network features as joint predictors. The performance metrics used in both works were adopted to evaluate the proposed probabilistic models as described in Section B.2.
@@
Most of the previously mentioned works use various approaches based on neural networks for NLP, but none of them has applied transformer models for the estimation of geolocation. The relatively new BERT model (released in 2018 [devlin2018bert]) is aimed at classification tasks and some works have already used it for country/city names predictions.
The base BERT model is intended to be finetuned on a downstream task which, in this case, was a regression to geographical coordinates (and GMM parameters such as weights and covariance) output as a type of sentence classification task. In general, BERT models have been pretrained using Masked Language Modeling (MLM) to improve the model’s understanding of bidirectional sentence representations, along with Next Sentence Prediction (NSP) to capture the inter-sentence relationships. Since BERT’s layers are hierarchical, early BERT layers learn more generic linguistic patterns, such as differences between multiple languages and their general sentence structures (in the case of the BERT multilingual model). While the later layers learn more task-specific patterns, in this case—the associations between points on the world map and specific terms referred to as LIWs, as well as text constructions and linguistic patterns commonly used in certain geographical areas.
@@
Nevertheless, the global dataset was multilingual so the corresponding BERT base model (’bert-base-multilingual-cased’) was chosen, which resulted in higher error distance metrics for the per-tweet and per-user geolocation prediction, as expected. A critical challenge addressed is the representation of multiple languages, since the ability to process multiple languages is a key factor in dividing text inputs into geospatial regions. To tackle this issue, a multilingual BERT model, pretrained on the 104 largest Wikipedia languages, was utilized as a starting point for finetuning on a global dataset with geolocation labels.
@@ 3 Materials and Methods
Step 2: importing default tokenizer and base BERT [devlin2018bert] model;
@@
The hyper-parameters tuning stage has covered such parameters as the type of the scheduler, minimal and maximal learning rates for the chosen scheduler, and the number of epochs. For the current task, the cosine type was chosen from the set of linear, cyclic, step, and plateau schedulers. Experiments have shown that the optimal learning rate reduction range starts at 1e-5 and ends at 1e-6 at the end of the last training epoch. The entering number of epochs was reduced from 5 to 3 for the dataset consisting of 2.7 million training samples, which are grouped into batches of 10 to 16 tweets in the model data loader. This decision was made as a result of the observed lack of significant error distance reduction after the third epoch for both training and test (development set) metrics. After each epoch, the models were evaluated on the development set, which consists of 300,000 samples. The described hyper-parameters were used for all proposed models, with variations only occurring in the loss function type, the number of outcomes (prediction points), covariance type, and the combination of text features.
@@ 3.1 Data preprocessing
Worth mentioning that the word-centric techniques, which use geocoders to recognize the referred place based on the geographic gazetteers [31]. The gazetteer typically includes entries for cities, towns, villages, landmarks, natural features, and other geographic entities, along with information such as their location, coordinates, elevation, etc. The challenge in this work was to find a substitute for the accurate knowledge base of geographical references (e.g., gazetteer terms).
@@
Considering the challenges of computational complexity and data availability, we chose to concentrate solely on user profile textual information (screen name, name, description, and location of the Twitter user profile), as shown in Table 1. This allowed us to effectively achieve our research objectives of predicting geolocation for the arbitrary tweets within our available resources and datasets. Worth mentioning, that the user’s home location wasn’t retrieved from the profile but rather sampled from the user’s tweets subset during evaluation.
@@ 4 Discussion and Conclusion
After the training performed on our Twitter archive data, the proposed model was able to predict the location of any Twitter user by processing the publicly available information posted online. While profile information provided by Twitter users is generally considered public, the combination or analysis of such data can have privacy implications. Tweets, being public by default, can also be used for analysis and publication. However, it’s worth noting that even seemingly innocuous information, when combined with other data points or analyzed in aggregate, can potentially reveal more about a person’s identity or habits. Therefore, while individual profile details may not be inherently sensitive, their combination or analysis might raise privacy concerns.
@@
As for the real-time user location monitoring, Yuan in the work [39] focused on tracking the movement of individuals or groups over time. Monitoring user geolocation in time can provide valuable insights for multiple applications such as disease outbreak trackers, urban mobility pattern analyzers, and location-based provider services. Such an analytical framework could be built on top of the proposed models, yet, in this study, only the commonly researched task of the user’s home location prediction was properly explored so far.
@@ 5 Acknowledgments
The authors acknowledge the Institute of Science and Technology (ISTA) for their material support and for granting access to the Twitter database archive, which was essential for the research.